OpenStudy (anonymous):

find f'(x) if the f(x) = the square root of 4 + e^2x

terenzreignz (terenzreignz):

Ever heard of chain rule? :)

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }\ln 4 + \frac{ 1 }{ 2}\ln e ^{2x}\]

OpenStudy (anonymous):

\[\large f(x)=(4+e^{2x})^{\frac{1}{2}}\] Can you use the chain rule like terenzreignz said above?

OpenStudy (anonymous):

i thought i did.\[(4+ e ^{2x} )^{1/2} = 1/2 \ln 4 + 1/2 \ln e ^{2x} = 1/2 (4x) + 1/2(2x) = 2x + x\]

terenzreignz (terenzreignz):

Any particular reason you put in a \(\ln\)?

OpenStudy (anonymous):

what else would i use?

OpenStudy (anonymous):

I'm pretty sure you do not need ln next to the 4.

OpenStudy (anonymous):

Nor the e^2x as well.

terenzreignz (terenzreignz):

hmm...\[\huge \frac{d}{dx}[f(x)]^n = n[f(x)]^{n-1}\cdot \frac{d}{dx}f(x)\]

OpenStudy (anonymous):

\[(4 + e ^{2x})^{3/2} \times 2x \times 2\]

OpenStudy (anonymous):

Are you differentiating or integrating?

OpenStudy (anonymous):

It seems you're getting confused with that.

OpenStudy (anonymous):

differentiating

OpenStudy (anonymous):

You're muddled up all over the place. First off, we're differentiating. And yes, you said it yourself just now.

OpenStudy (anonymous):

So if we're differentiating using the chain rule, we first move the exponent to the front of the expression.

OpenStudy (anonymous):

Then we differentiate the inside. After that, we rewrite the original expression but with the exponent being less than 1 of the original exponent.

OpenStudy (anonymous):

It's very simple if you take everything Step by STEP. I put emphasis on step by step because that's where people get stumped on. They have a need to skip steps.

OpenStudy (anonymous):

That's where most of your silly mistakes come from. skipping steps.

OpenStudy (anonymous):

i thought in order to bring the exponent to the front i needed a ln

OpenStudy (anonymous):

\[1/2 (4 + e ^{2x}) = 1/2 ( 0 + e^{2x-1}) = 1/2 e^{2x-1}\]

OpenStudy (anonymous):

Ah. I'm not sure how you thought of that, but I will help you clear everything and say that ln isn't needed in here. ;)

OpenStudy (anonymous):

Why is e^{2x-1} in there. I feel that you though e stood for exponent.

OpenStudy (anonymous):

The exponent is right beside the brackets/paranthesis.

OpenStudy (anonymous):

thought*

OpenStudy (anonymous):

and you differentiate e^2x. I think I told you this already about differentiating e. You rewrite the whole thing again and then you differentiate 2x.

OpenStudy (anonymous):

the 2x comes from the exponent of e. Now when we use chain rule for the whole expression, you change the exponent of the whole.

OpenStudy (anonymous):

the whole meaning the exponent right at the very end which was 1/2. What's 1/2-1?

OpenStudy (anonymous):

the "whole", *

OpenStudy (anonymous):

no .. im trying to work so that i use all of the rules you told me so what am i doing wrong?

OpenStudy (anonymous):

Okay when I say differentiate the inside, what method of differentiating would you use?

OpenStudy (anonymous):

You're using the chain rule as a WHOLE. Am I right? And in the WHOLE, you have a certain part where you differentiate the inside.

OpenStudy (anonymous):

that certain part is where you differentiate normally.

OpenStudy (anonymous):

That's where I think you're getting muddle up. You think when you differentiate the inside, you're also using the chain rule to differentiate.

OpenStudy (anonymous):

Logically speaking if you were to differentiate the inside and you're differentiating e^2x, what makes yout hink you need to use the chain rule to differentiate it? hmmm...seems a bit strange, don't you think? ;)

OpenStudy (anonymous):

you think*

OpenStudy (anonymous):

im more confused than when i started.

OpenStudy (anonymous):

\[\large f(x)=(4+e^{2x})^{\frac{1}{2}}\] Okay. we have this right. the square root sign can be switched into a power/exponent which is 1/2.

OpenStudy (anonymous):

we have this right?*

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now we use the chain rule. It seems you're not really familiar with the chain rule.

OpenStudy (anonymous):

This certain area I think you need more practice on. Not trying to lower your standards of thinking, but it is a good thing to practice on the critical parts of maths such as this.

OpenStudy (anonymous):

Okay, the chain rule: Let's start by moving the power/exponent to the front of the expression first.We're going step by step so you don't get confused as much. \[\large f'(x)=\frac{1}{2}(4+e^{2x})^{\frac{1}{2}}\]

OpenStudy (anonymous):

You confused about this first step or not?

OpenStudy (anonymous):

I juse placed 1/2 to the front. That's your very first step.

OpenStudy (anonymous):

if i pulled it to the front then why is it still in the exponent place?

OpenStudy (anonymous):

I'm doing this stpe by step. How am I going to do this step by step. I have no other way because I'm on the internet. If I was teaching you in person, I would be able to teach you stpe by step. YOu just have to bare with me on this.

OpenStudy (anonymous):

\[\large f'(x)=\frac{1}{2}(0+2e^{2x})\] Then we differentiate the inside.

OpenStudy (anonymous):

After that, we rewrite the original expression. THe power will become one less than before. \[\large f'(x)=\frac{1}{2}(0+2e^{2x})(4+e^{2x})^{\frac{1}{2}-1}\]

OpenStudy (anonymous):

Then you can get rid of the 0 and simplify that. And that will be your final answer.

OpenStudy (anonymous):

okay .. i know what i was missing.i didn't multiply it by the original equ.

OpenStudy (anonymous):

It's simple if you don't Over-analyse things. It's very easy to over-analyse. Okay, so you understand everything I said or are you still confused and all?

OpenStudy (anonymous):

no i got got it.\[\frac{ e ^{2x} }{ \sqrt{4 + e ^{2x}}}\]

OpenStudy (anonymous):

Yep. Well done. Hope you don't make the same mistake when using the chain rule again. I myself would be disappointed if I made the same mistake over and over again.

OpenStudy (anonymous):

i get it now! thank you .. got all the other ones correct. but how would i find the derivative of \[xe ^{y} + 1 = xy\]\[\frac{ xe ^{y} + 1 }{ x} = y\]\[xe ^{y} + 1 \times x ^{-1}\] is that correct so far?

OpenStudy (anonymous):

differentiate with respect to x?

OpenStudy (anonymous):

It seems that you can implicitly differentiate. And that's not correct.

OpenStudy (anonymous):

When you differentiate xe^y, you must use the product rule. Same goes with differentiating xy.

OpenStudy (anonymous):

Your differentiation skills in general out a bit rusty and out of touch. But the more you practice, the better you get. The only thing I stress on would be practicing on your basics of differentiation. For example, practice on the chain rule, product rule, quotient rule, etc. All of these methods of differentiation should be considered when practicing.

OpenStudy (anonymous):

general are*

OpenStudy (anonymous):

first times the the dx/dy of the second plus second times the dx/dy of the first